# Cellular, Cell-less, and Everything in Between: A Unified Framework for Utility Region Analysis in Wireless Networks

### v2

@@ II-A Notation and standard definitions
Consider the general interference mapping $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto{\bm{Mx}}$, where ${\bm{M}}\in{\mathbb{R}}_{++}^{N\times N}$ is an inverse Z-matrix. Then $G$ is norm inducing, and the induced norm satisfies $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~\|{\bm{x}}\|_{G}=\rho({\mathrm{diag}(\bm{x})}{\bm{M}})$.
The next example, which is also relevant to the wireless systems discussed in the next section, shows a norm-inducing general interference mapping $T_{\|\cdot\|}$} that is constructed from an explicitly given mapping $T$ satisfying hypothesis ($T{T_{\|\cdot\|}}$) and a monotone norm. Moreover, for every ${\bm{x}}\in{\mathbb{R}}_{+}^{N}$, the norm $\|{\bm{x}}\|_{T_{\|\cdot\|}}$ can be easily computed from the spectral radii of matrices. The proof of the statement in the next example is shown in the Supplementary Material.
@@
III System model
We now focus on the SINR as the utility function for the uplink of a wireless system with $N\in\mathbb{N}$ single-antenna users, which are components of the set $\mathcal{N}:=\{1,\ldots,N\}$. However, many results we discuss below can also be extended to the downlink using duality arguments [36, 37]. To avoid technology specific models, we refrain from specifying the number of access points, the number of antennas at these access points, or the particular wireless architecture in use (e.g., cellular massive MIMO, cell-less, etc.). For many of these technologies, given the uplink power allocation ${\bm{p}}=(p_{1},\ldots,p_{N})\in{\mathbb{R}}_{++}^{N}$, the SINR $s_{n}$ for each user $n\in\mathcal{N}$ can be expressed in a unified form as follows:
$(\forall n\in\mathcal{N})~s_{n}=\sup_{{\bm{y}}\in\mathcal{Y}_{n}}\dfrac{b_{n}({\bm{y}})~p_{n}}{{\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}+\sigma_{n}({\bm{y}})}$ (15)
where $\mathcal{Y}_{n}$ is typically a set of random or deterministic vectors and, for every ${\bm{y}}\in\mathcal{Y}_{n}$, the parameters $b_{n}({\bm{y}})\in{\mathbb{R}}_{++}$, ${\bm{c}}_{n}({\bm{y}})={\mathbb{R}}_{+}^{K}$, and $\sigma_{n}({\bm{y}})\in{\mathbb{R}}_{++}^{N}$ are system specific – see Example 3 below. We assume that, for each $n\in\mathcal{N}$, the function $\mathcal{Y}_{n}\to{\mathbb{R}}_{++}:{\bm{y}}\mapsto\sigma_{n}({\bm{y}})$ is bounded away from zero (i.e., $(\exists\delta>0)(\forall{\bm{y}}\in\mathcal{Y}_{n})(\forall n\in\mathcal{N})~\sigma_{n}({\bm{y}})\geq\delta$), which is a mild assumption that typically holds in networks with noise. To avoid technical digressions of little practical relevance, we also assume that the supremum in (15) is always attained and, for every $n\in\mathcal{N}$, we have $s_{n}\in{\mathbb{R}}_{++}$ if ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$. In most current studies, $\mathcal{Y}_{n}$ is the set of available beamformers for user $n\in\mathcal{N}$, and the remaining parameters, which are functions of the beamformers, are constructed based on information about the channels, thermal noise, and the quality of channel estimates. Particular cases of this model can be found in, for instance, cell-less networks [6, Ch. 3, Ch. 7][4, 21, 15, 5], massive MIMO systems [7, Ch.5.3.2], [8, Theorem7.1], and conventional cellular networks [13, Ch. 4].
@@ III System model
(Two-time scale beamforming in cell-less systems [21, 4, 5]) In the uplink of a cell-less network with single-antenna users, let the channel of user $n\in\mathcal{N}$ across all $M\in{\mathbb{N}}$ access points, each equipped with $L\in{\mathbb{N}}$ antennas, be a $\mathbb{C}^{ML}$-valued random vector ${\bm{h}}_{n}$. Assume a fixed beamforming strategy; i.e., for each user $n\in\mathcal{N}$, the set $\mathcal{Y}_{n}$ is a singleton containing a $\mathbb{C}^{ML}$-valued random vector ${\bm{v}}_{n}$. For instance, if the access points use as their local beamformers the traditional maximum ratio combining strategy, the random vector ${\bm{v}}_{n}$ can be interpreted as an estimate of the instantaneous channel ${\bm{h}}_{n}$. Nevertheless, we emphasize that other strategies (such as zero-forcing beamforming) are allowed, as long as the beamformer remains unchanged if the power allocation $\bm{p}$} changes. This assumption is adopted in [4, 5] and also in the numerous references cited therein. To compute the well-known use-and-then-forget (UatF) ergodic capacity inner bound [7] for user $n\in{\mathbb{N}}$, we use the following SINR expression for a given power allocation ${\bm{p}}=(p_{1},\ldots,p_{N})$:
$s_{n}=\dfrac{p_{n}|E[{\bm{h}}_{n}^{H}{\bm{v}}_{n}]|^{2}}{p_{n}V({\bm{h}}^{H}_{n}{\bm{v}}_{n})+\sum\limits_{k\in\mathcal{N}\backslash\{n\}}p_{k}~E[|{\bm{h}}_{k}^{H}{\bm{v}}_{n}|^{2}]+E[\|{\bm{v}}_{n}\|_{2}^{2}]},$ (16)
@@
An essential component to take note in this model is the self-interference term $V({\bm{h}}^{H}_{1}{\bm{v}}_{1})$ (which is also related to the notion of beamforming gain uncertainty in the massive MIMO literature [7]). This term is crucial in our analysis of the shape of the SINR and achievable rate regions, as explained later in Sect. IV-A.
We can easily establish a connection between (15) and the discussion in Sect. I-A. More specifically, the function $t_{n}$ in (1) for user $n\in\mathcal{N}$ is given by
@@
and $s_{n}$ in (15) is the value taken by the utility $u_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ in (1) of user $n\in\mathcal{N}$ for a given power vector ${\bm{p}}\in{\mathbb{R}}^{N}_{++}$; i.e., $u_{n}({\bm{p}})=p_{n}/t_{n}({\bm{p}})=s_{n}$. The function $t_{n}$ in (17) is a standard interference function for every $n\in\mathcal{N}$ because, for each ${\bm{y}}\in\mathcal{Y}_{n}$, the function ${\bm{p}}\mapsto({\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}+\sigma_{n}({\bm{y}}))/b_{n}({\bm{y}})$ is affine (hence concave), so $t_{n}$ is concave because it is the pointwise infimum of concave functions [38, Proposition 8.16]. Furthermore, $t_{n}$ is positive everywhere in the nonnegative cone ${\mathbb{R}}_{+}^{N}$ by assumption (which in particular implies that $(\forall n\in\mathcal{N})\inf_{{\bm{y}}\in\mathcal{Y}_{n}}\sigma_{n}({\bm{y}})/b_{n}({\bm{y}})>0$), so, being positive and concave, $t_{n}$ is a standard interference function [39, Proposition 1].
@@
where $\|\cdot\|$ is a monotone norm. Power constraints of this type are fairly general because any set with nonempty interior that is compact, convex, and downward comprehensive in ${\mathbb{R}}_{+}^{N}$ can be expressed as in (18) for some monotone norm $\|\cdot\|$ [35, Proposition 2]. Readers may notice that set $\mathcal{P}$ in (18) differ from the set $\mathcal{P}$ already introduced in Sect. I-A because the scalar $p_{\mathrm{max}}\in{\mathbb{R}}_{++}$ is omitted in (18). However, this omission does not result in any loss of generality. Given any monotone norm $\|\cdot\|_{\star}$, the constraint $\|{\bm{x}}\|_{\star}\leq p_{\mathrm{max}}$ can be equivalently expressed as the set $\mathcal{P}$ in (18) by considering the scaled monotone norm $(\forall{\bm{x}}\in{\mathbb{R}}^{N})~\|{\bm{x}}\|:=(1/p_{\mathrm{max}})\|{\bm{x}}\|_{\star}$, which satisfies $(\forall{\bm{x}}\in{\mathbb{R}}^{N})~\|{\bm{x}}\|\leq 1\Leftrightarrow\|{\bm{x}}\|_{\star}\leq{p_{\mathrm{max}}}$.
To rewrite (15) in a form suitable for analysis via fixed-point theory, given a power vector ${\bm{p}}\in{\mathbb{R}}^{N}_{++}$, we isolate each coordinate of this power vector and use the coordinate functions defined in (17) to obtain $(\forall n\in\mathcal{N})~p_{n}=s_{n}t_{n}({\bm{p}}),$ which, in vector form, can be equivalently expressed as
${\bm{p}}={\mathrm{diag}(\bm{s})}T({\bm{p}}),$ (19)
where $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto\left[t_{1}({\bm{x}}),\cdots,t_{N}({\bm{x}})\right]^{t}$ and ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$ is the vector of SINR values.
To relate the above model to results derived in Sect. II-B, note that $T$ in (19) is a mapping satisfying hypothesis ($T,{T_{\|\cdot\|}}$) in Definition 4. Therefore, considering the monotone norm $\|\cdot\|$ in the definition of the set $\mathcal{P}$ in (18), we deduce that the corresponding mapping $T_{\|\cdot\|}$} in Definition 4 is given by
@@
${(t_{n})_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}:{\bm{p}}\mapsto\inf_{{\bm{y}}\in\mathcal{Y}_{n}}\left(\dfrac{{\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}}{b_{n}({\bm{y}})}+\dfrac{\sigma_{n}({\bm{y}})}{b_{n}({\bm{y}})}\|{\bm{p}}\|\right).$
III-A Achievable rate and SINR regions in the absence of power constraints
To study the feasible SINR and rate regions of the wireless model introduced above, let us start by ignoring the presence of the power constraint $\mathcal{P}$ to relax key assumptions made in [13, Ch. 5], particularly the linearity of the interference models and the symmetry of the interference matrices (cf. [13, Ch. 5.4.4]).
The equality in (19) proves that, if the SINR requirements described by the positive vector $\bm{s}$} are achievable by the power vector ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$, then
${\bm{p}}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T):=\{{\bm{p}}\in{\mathbb{R}}_{++}^{N}~|~{\mathrm{diag}(\bm{s})}T({\bm{p}})={\bm{p}}\},$ (21)
where $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ is the standard interference mapping defined above. Therefore, the set $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$ is either a singleton or the empty set [22], and we verify that the SINR requirements represented by vector $\bm{s}$} are achievable by a power vector ${\bm{p}}\in{\mathbb{R}}_{++}^{N}$ if and only if $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$, and this power vector is unique if it exists [22]. In compact form, the set of positive SINR values that the network can provide to the users can be expressed as $\mathcal{S}=\{{\bm{s}}\in{\mathbb{R}}_{++}^{N}\mid\mathrm{Fix}(\mathrm{diag}({\bm{s}})T)\neq\emptyset\}.$ This approach based on fixed point theory has been introduced to the wireless literature in [22], and it has since seen significant advancements in both theory and practice [1, 18, 23, 19, 37].
In the absence of power constraints, we can use the concept of asymptotic mappings (Definition 2) to obtain a second characterization of the set $\mathcal{S}$ of positive, achievable SINR levels ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$. More precisely, given ${\bm{s}}\in{\mathbb{R}}^{N}_{++}$, continuity of the mapping ${\bm{u}}\mapsto\mathrm{diag}({\bm{s}}){\bm{u}}$ yields
$(\forall{\bm{s}}\in{\mathbb{R}}_{++}^{N})(\forall{\bm{x}}\in{\mathbb{R}}^{N})\\ (\mathrm{diag}({\bm{s}})T)_{\infty}({\bm{x}})=\lim_{h\to\infty}(1/h)\mathrm{diag}({\bm{s}})T(h{\bm{x}})\\ =\mathrm{diag}({\bm{s}})\lim_{h\to\infty}(1/h)T(h{\bm{x}})=\mathrm{diag}({\bm{s}})T_{\infty}({\bm{x}}),$
which, together with [23, Proposition 4] (Fact 9 in the Supplemental Material), leads to the following alternative characterization of $\mathcal{S}$:

### v3

@@ II-A Notation and standard definitions
Consider the general interference mapping $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto{\bm{Mx}}$, where ${\bm{M}}\in{\mathbb{R}}_{++}^{N\times N}$ is an inverse Z-matrix. Then $G$ is norm-inducing, and the induced norm satisfies $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~\|{\bm{x}}\|_{G}=\rho({\mathrm{diag}(\bm{x})}{\bm{M}})$.
Proof.
Standard Perron-Frobenius theory shows that the spectral radius $\rho(G)$ is positive, and there exists a positive left eigenvector ${\bm{l}}\in{\mathbb{R}}^{N}_{++}$ and a positive right eigenvector ${\bm{r}}\in{\mathbb{R}}_{++}^{N}$ associated with $\rho(G)$. Choose $\epsilon>0$ such that ${\bm{M}}_{1}={\bm{M}}-\epsilon~{\bm{lr}}^{t}\in{\mathbb{R}}_{+}^{N\times N}$ (a scalar $\epsilon$ with this property always exists because $\bm{M}$} is positive). The mapping $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{x}}\mapsto{\bm{M}}_{1}{\bm{x}}+\epsilon{\bm{l}}$ is a standard interference mapping because it is positive and concave [11, Proposition 1]. Using the monotone norm $(\forall{\bm{x}}\in{\mathbb{R}}^{N})~\|{\bm{x}}\|_{\bm{r}}:={\bm{r}}^{t}|{\bm{x}}|$, we have $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})~T_{\|\cdot\|_{\bm{r}}}({\bm{x}})=G({\bm{x}})={\bm{Mx}}={\bm{M}}_{1}{\bm{x}}+\epsilon{\bm{l}}\|{\bm{x}}\|_{\bm{r}}$, so $T$ and $T_{\|\cdot\|_{\bm{r}}}$ satisfy hypothesis $(T,T_{\|\cdot\|_{\bm{r}}})$. Therefore, we are in the setting of Definition 5 because $g_{G}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}^{N}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}})G)=\rho(\mathrm{diag}({\bm{x}}){\bm{M}})$ is convex on ${\mathbb{R}}_{+}^{N}$ as a direct consequence of Fact 5 in the Supplemental Material, so $G$ is norm-inducing as claimed, and the induced norm satisfies $(\forall{\bm{x}}\in{\mathbb{R}}_{+}^{N})\|{\bm{x}}\|_{G}=\rho({\mathrm{diag}(\bm{x})}{\bm{M}})$. ∎
Remark 2.
The function $G:{\mathbb{R}}^{N}_{+}\to{\mathbb{R}}_{+}:{\bm{x}}\mapsto\rho(\mathrm{diag}({\bm{x}}){\bm{M}})$ is also convex if ${\bm{M}}\in{\mathbb{R}}_{+}^{N\times N}$ is either (i) irreducible and symmetric positive semidefinite [35, Theorem 1.51], or (ii) triangular (upper or lower). In the triangular case, for a given matrix ${\bm{M}}\in{\mathbb{R}}^{N\times N}_{+}$, the function $G:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}:{\bm{x}}=(x_{1},\dots,x_{N})\mapsto\rho(\mathrm{diag}({\bm{x}}){\bm{M}})=\max_{n\in\{1,\ldots,N\}}x_{n}{\bm{M}}_{n,n}$ (NOTE: ${\bm{M}}_{n,n}$ denotes the $n$th diagonal entry of $\bm{M}$}) is the maximum of convex (linear) functions, hence convex. Note that nonnegative matrices in the classes above need not be inverse Z-matrices or norm-inducing. We do not pursue these cases here, but they may be useful for establishing convexity of utility regions beyond the scope of the present study.
The next result establishes a sufficient condition ensuring that a mapping $T_{\|\cdot\|}$} of the type considered in Example 2 is norm-inducing.
@@
Proof.
[…]
III Applications to wireless networks
III-A Interference models
We now focus on the SINR as the utility function for the uplink of a wireless system with $N\in\mathbb{N}$ single-antenna users, which are components of the set $\mathcal{N}:=\{1,\ldots,N\}$. However, some results we discuss below can also be extended to the downlink using duality arguments [3, 34, 24][2, Sect. 4.3.2]. To avoid technology specific models, we refrain from specifying the number of access points, the number of antennas at these access points, or the particular wireless architecture in use (e.g., cellular massive MIMO, cell-less, etc.). For many of these technologies, given the uplink power allocation ${\bm{p}}^{\star}=(p_{1}^{\star},\ldots,p_{N}^{\star})\in{\mathbb{R}}_{++}^{N}$, the SINR $s_{n}$ for each user $n\in\mathcal{N}$ can be expressed in a unified form as follows:
$(\forall n\in\mathcal{N})~s_{n}=\sup_{{\bm{y}}\in\mathcal{Y}_{n}}\dfrac{b_{n}({\bm{y}})~p_{n}^{\star}}{{\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}^{\star}+\sigma_{n}({\bm{y}})},$ (20)
where $\mathcal{Y}_{n}$ is typically a set of random or deterministic vectors and, for every ${\bm{y}}\in\mathcal{Y}_{n}$, the parameters $b_{n}({\bm{y}})\in{\mathbb{R}}_{++}$, ${\bm{c}}_{n}({\bm{y}})={\mathbb{R}}_{+}^{K}$, and $\sigma_{n}({\bm{y}})\in{\mathbb{R}}_{++}^{N}$ are system specific – see Examples 6 and 7 below. We assume that, for each $n\in\mathcal{N}$, the function $\mathcal{Y}_{n}\to{\mathbb{R}}_{++}:{\bm{y}}\mapsto\sigma_{n}({\bm{y}})$ is bounded away from zero (i.e., $(\exists\delta>0)(\forall n\in\mathcal{N})(\forall{\bm{y}}\in\mathcal{Y}_{n})~\sigma_{n}({\bm{y}})\geq\delta$), which is a mild assumption that typically holds in networks with noise. To avoid technical digressions of little practical relevance, we also assume that the supremum in (20) is always attained and, for every $n\in\mathcal{N}$, we have $s_{n}\in{\mathbb{R}}_{++}$ if ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$. As Example 6 below shows, in current studies, $\mathcal{Y}_{n}$ is often (though not always – see Example 7) the set of available beamformers for user $n\in\mathcal{N}$, and the remaining parameters, which are functions of the beamformers, are constructed based on information about the channels, thermal noise, and the quality of channel estimates. Particular cases of this model can be found in, for instance, cell-less networks [15, Ch. 3, Ch. 7][26, 27, 34, 25, 12], massive MIMO systems [23, Ch.5.3.2], [2, Theorem7.1], and conventional cellular networks [35, Ch. 4]. The next two examples illustrate how to specialize (20) to model modern cell-less networks and classical cellular networks with optimal cell-site selection.
@@ III-A Interference models
(Two-time scale receive beamforming in cell-less systems [27, 26, 12]) In the uplink of a cell-less network with single-antenna users, let the channel of user $n\in\mathcal{N}$ across all $M\in{\mathbb{N}}$ access points, each equipped with $L\in{\mathbb{N}}$ antennas, be a $\mathbb{C}^{ML}$-valued random vector ${\bm{h}}_{n}$. Assume a fixed (digital) receive beamforming strategy; i.e., for each user $n\in\mathcal{N}$, the set $\mathcal{Y}_{n}$ is a singleton containing a $\mathbb{C}^{ML}$-valued random vector ${\bm{v}}_{n}$. For instance, if the access points use as their local receive beamformers the traditional maximum ratio combining strategy, the random vector ${\bm{v}}_{n}$ can be interpreted as an estimate of the instantaneous channel ${\bm{h}}_{n}$. Nevertheless, we emphasize that other strategies (such as zero-forcing receive beamforming) are allowed, as long as the receive beamformer remains unchanged if the power allocation changes. This assumption is adopted in [26, 12] and also in the numerous references cited therein. To compute the well-known use-and-then-forget (UatF) ergodic capacity inner bound [23] for user $n\in\mathcal{N}$, we use the following SINR expression for a given power allocation ${\bm{p}}^{\star}=(p_{1}^{\star},\ldots,p_{N}^{\star})\in{\mathbb{R}}_{++}^{N}$:
$s_{n}=\dfrac{p^{\star}_{n}|E[{\bm{h}}_{n}^{H}{\bm{v}}_{n}]|^{2}}{p^{\star}_{n}V({\bm{h}}^{H}_{n}{\bm{v}}_{n})+\sum\limits_{k\in\mathcal{N}\backslash\{n\}}p^{\star}_{k}~E[|{\bm{h}}_{k}^{H}{\bm{v}}_{n}|^{2}]+E[\|{\bm{v}}_{n}\|_{2}^{2}]},$ (21)
@@
An essential component to note in this model is the self-interference term $V({\bm{h}}^{H}_{1}{\bm{v}}_{1})$ (which is also related to the notion of beamforming gain uncertainty in the massive MIMO literature [23]). This term is crucial in our analysis of the shape of the SINR and achievable rate regions, as explained later in Sect. IV-A.
Example 7.
(Cellular systems with optimal cell-site selection [18]) Consider the uplink of a cellular network with $Y\in{\mathbb{N}}$ base stations indexed by the set $\mathcal{Y}=\{1,\ldots,Y\}$, and let the $N\in{\mathbb{N}}$ users be indexed by the set $\mathcal{N}=\{1,\ldots,N\}$. Let $b_{n}(y)\in{\mathbb{R}}_{++}$ denote the channel gain between user $n\in\mathcal{N}$ and base station $y\in\mathcal{Y}_{n}$, where $\mathcal{Y}_{n}\subset\mathcal{Y}$ represents the set of base stations to which user $n\in\mathcal{N}$ is allowed to connect. For simplicity, assume that the noise level $\sigma\in{\mathbb{R}}_{++}$ is identical for all links.
Define the interference channel vector for user $n\in\mathcal{N}$ at base station $y\in\mathcal{Y}_{n}$ by ${\bm{c}}_{n}(y)=[c_{1,n}(y),\cdots,c_{N,n}(y)]^{t}\in{\mathbb{R}}_{++}^{N}$, where $c_{k,n}(y)=b_{k}(y)$ if $k\neq n$ and 0 otherwise. If ${\bm{p}}^{\star}=[p^{\star}_{1},\ldots,p^{\star}_{N}]^{t}\in{\mathbb{R}}_{++}^{N}$ is the uplink transmit power vector for all users, and each user $n\in\mathcal{N}$ connects to its best serving base station, then the resulting SINR of user $n\in\mathcal{N}$ is given by
$s_{n}=\max_{y\in\mathcal{Y}_{n}}\dfrac{b_{n}(y)p_{n}^{\star}}{{\bm{c}}_{n}(y)^{t}{\bm{p}}^{\star}+\sigma},$
which is a particular case of (20).
In the remainder of this study, Example 6 will be used extensively to illustrate the main implications of our results. Nevertheless, a few remarks regarding channel acquisition are in order.
Remark 3.
As noted in Example 6, the beamformers in (21) may be designed based on channel estimates rather than the true channels, so we naturally account for channel estimation errors. In the sequel, however, we assume that the channel estimation quality – and the resources allocated to channel estimation – are independent of the particular set of users being served. Relaxing this assumption would substantially complicate the analysis. In particular, regardless of whether the achievable-rate region is convex or nonconvex, determining whether users benefit from being grouped would then require an explicit model of the estimation procedure (e.g., how pilot overhead and the resulting estimation accuracy change when users are partitioned into smaller or larger groups).
We can also establish a connection between (20) and the discussion in Sect. I-A. More specifically, the function $t_{n}$ in (1) for user $n\in\mathcal{N}$ is given by
@@
(which we assume to be continuous) and $s_{n}$ in (20) is the value taken by the utility $U_{n}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{+}$ in (1) of user $n\in\mathcal{N}$ for a given power vector ${\bm{p}}^{\star}\in{\mathbb{R}}^{N}_{++}$; i.e., $U_{n}({\bm{p}}^{\star})=p^{\star}_{n}/t_{n}({\bm{p}}^{\star})=s_{n}$. The function $t_{n}$ in (22) is a standard interference function for every $n\in\mathcal{N}$ because, for each ${\bm{y}}\in\mathcal{Y}_{n}$, the function ${\bm{p}}\mapsto({\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}+\sigma_{n}({\bm{y}}))/b_{n}({\bm{y}})$ is affine (hence concave), so $t_{n}$ is concave because it is the pointwise infimum of concave functions [1, Proposition 8.16]. [^4] For every $n\in\mathcal{N}$, the restriction $t_{n}:{\mathbb{R}}^{N}_{++}\to{\mathbb{R}}_{+}$ of $t_{n}$ to the domain ${\mathbb{R}}_{++}^{N}$ admits a continuous, concavity-preserving extension to the domain ${\mathbb{R}}_{+}^{N}$ [32, Theorem 10.3] that is also bounded away from zero. Therefore, we can always redefine $t_{n}$ to be this continuous extension and assume continuity of $t_{n}$ on ${\mathbb{R}}_{+}^{N}$. Furthermore, $t_{n}$ is positive everywhere on the nonnegative cone ${\mathbb{R}}_{+}^{N}$ by assumption (which in particular implies that $(\forall n\in\mathcal{N})\inf_{{\bm{y}}\in\mathcal{Y}_{n}}\sigma_{n}({\bm{y}})/b_{n}({\bm{y}})>0$), so, being positive and concave, $t_{n}$ is a standard interference function [11, Proposition 1].
@@
where $\|\cdot\|$ is a monotone norm. Power constraints of this type are fairly general because any set with nonempty interior that is compact, convex, and downward comprehensive in ${\mathbb{R}}_{+}^{N}$ can be expressed as in (23) for some monotone norm $\|\cdot\|$ [8, Proposition 2]. Readers may notice that set $\mathcal{P}$ in (23) differs from the set $\mathcal{P}$ already introduced in Sect. I-A because the scalar $p_{\mathrm{max}}\in{\mathbb{R}}_{++}$ is omitted in (23). However, this omission does not result in any loss of generality. Given any monotone norm $\|\cdot\|_{\star}$, the constraint $\|{\bm{p}}\|_{\star}\leq p_{\mathrm{max}}$ can be equivalently expressed as the set $\mathcal{P}$ in (23) by considering the scaled monotone norm $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|:=(1/p_{\mathrm{max}})\|{\bm{p}}\|_{\star}$, which satisfies $(\forall{\bm{p}}\in{\mathbb{R}}^{N})~\|{\bm{p}}\|\leq 1\Leftrightarrow\|{\bm{p}}\|_{\star}\leq{p_{\mathrm{max}}}$.
To rewrite (20) in a form suitable for analysis via fixed point theory, we isolate $p_{n}^{\star}\in{\mathbb{R}}_{++}$ in the numerator of the fraction inside the supremum for each $n\in\mathcal{N}$. Using the coordinate functions defined in (22), we obtain $(\forall n\in\mathcal{N})~p_{n}^{\star}=s_{n}t_{n}({\bm{p}}^{\star}),$ which, in vector form, can be equivalently expressed as
${\bm{p}}^{\star}={\mathrm{diag}(\bm{s})}T({\bm{p}}^{\star}),$ (24)
where $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}:{\bm{p}}\mapsto\left[t_{1}({\bm{p}}),\cdots,t_{N}({\bm{p}})\right]^{t}$ and ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$ is the vector of SINR values.
To relate the above model to the results derived in Sect. II-B – and, in particular, Definition 4 – consider the monotone norm $\|\cdot\|$ in the definition of the set $\mathcal{P}$ in (23), and define
@@
${(t_{n})_{\|\cdot\|}}:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}:{\bm{p}}\mapsto\inf_{{\bm{y}}\in\mathcal{Y}_{n}}\left(\dfrac{{\bm{c}}_{n}({\bm{y}})^{t}{\bm{p}}}{b_{n}({\bm{y}})}+\dfrac{\sigma_{n}({\bm{y}})}{b_{n}({\bm{y}})}\|{\bm{p}}\|\right)$
(which we also assume to be continuous). We verify that $T$ and $T_{\|\cdot\|}$} defined as above satisfy hypothesis $(T,{T_{\|\cdot\|}})$ for the monotone norm $\|\cdot\|$ used in (23).
III-B Achievable rate and SINR regions in the absence of power constraints
To study the feasible SINR and rate regions of the wireless model introduced above, let us start by ignoring the presence of the power constraint $\mathcal{P}$ to relax key assumptions made in [35, Ch. 5], particularly the linearity of the interference models and the symmetry of the interference matrices (cf. [35, Ch. 5.4.4]).
The equality in (24) proves that, if the SINR requirements described by the positive vector $\bm{s}$} are achievable by the power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$, then
${\bm{p}}^{\star}\in\mathrm{Fix}({\mathrm{diag}(\bm{s})}T):=\{{\bm{p}}\in{\mathbb{R}}_{++}^{N}~|~{\mathrm{diag}(\bm{s})}T({\bm{p}})={\bm{p}}\},$ (26)
where $T:{\mathbb{R}}_{+}^{N}\to{\mathbb{R}}_{++}^{N}$ is the standard interference mapping defined below (24). Therefore, the set $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)$ is either a singleton or the empty set [42], and we verify that the SINR requirements represented by the vector $\bm{s}$} are achievable by a power vector ${\bm{p}}^{\star}\in{\mathbb{R}}_{++}^{N}$ if and only if $\mathrm{Fix}({\mathrm{diag}(\bm{s})}T)\neq\emptyset$, and this power vector is unique if it exists [42]. In compact form, the set of positive SINR values that the network can provide to the users can be expressed as $\mathcal{S}=\{{\bm{s}}\in{\mathbb{R}}_{++}^{N}\mid\mathrm{Fix}(\mathrm{diag}({\bm{s}})T)\neq\emptyset\}.$ This approach based on fixed point theory has been introduced to the wireless literature in [42], and it has since seen significant advancements in both theory and practice [29, 33, 9, 31, 24].
In the absence of power constraints, we can use the concept of asymptotic mappings (Definition 2) to obtain a second characterization of the set $\mathcal{S}$ of positive, achievable SINR levels ${\bm{s}}=(s_{1},\ldots,s_{N})\in{\mathbb{R}}_{++}^{N}$. More precisely, given ${\bm{s}}\in{\mathbb{R}}^{N}_{++}$, continuity of the mapping ${\bm{x}}\mapsto\mathrm{diag}({\bm{s}}){\bm{x}}$ yields
$(\forall{\bm{s}}\in{\mathbb{R}}_{++}^{N})(\forall{\bm{p}}\in{\mathbb{R}}^{N})\\ (\mathrm{diag}({\bm{s}})T)_{\infty}({\bm{p}})=\lim_{h\to\infty}(1/h)\mathrm{diag}({\bm{s}})T(h{\bm{p}})\\ =\mathrm{diag}({\bm{s}})\lim_{h\to\infty}(1/h)T(h{\bm{p}})=\mathrm{diag}({\bm{s}})T_{\infty}({\bm{p}}),$
which, together with Fact 9 in the Supplemental Material, leads to the following alternative characterization of $\mathcal{S}$: